A particle of mass 1 kg starts moving so that its speed varies as b√s where b=1 unit and S is distance covered by the particle in time T find kinetic energy of particle at t=1 sec.

Grade:11

1 Answers

anusree
5 years ago
Speed=v=b√s
V=s/T
S=vT
V=b√vT
V/√v=b√T
√v=b√T
Square on both sides
V=BT
b=1unit
T=t=1s
So,v=1
K.E=1/2mv^2
m=1kg
V=1m/s
So K.E =1/2=0.5j